OpenStudy (anonymous):

The density of water is 1.00 h/mL at 4oC. How many water molecules are present in 2.46 mL of water at this temperature? I'm really confused how to figure this out.

OpenStudy (somy):

what is the formula of density?

OpenStudy (anonymous):

d=m/v

OpenStudy (somy):

good so find m

OpenStudy (anonymous):

1.00= m/2.46 so you would divide 1.00 by 2.46...right?

OpenStudy (somy):

nope \[\sf \frac{ D }{ 1 } = \frac{ m }{ V }\] cross multiplication same as in math

OpenStudy (somy):

D is same as D/1

OpenStudy (anonymous):

So 1.00*2.46 then?

OpenStudy (anonymous):

giving 2.46?

OpenStudy (somy):

yep

OpenStudy (somy):

so that is your mass in grams now

OpenStudy (somy):

now you know mass and you know that talk is about H2O which means you can calculate molecular mass of H2O using periodic table

OpenStudy (somy):

thus you can find mole

OpenStudy (somy):

you know the formula of mole?

OpenStudy (anonymous):

You times the number of the element you have by the mass. So H (2*1.008) = 2.016 right?

OpenStudy (anonymous):

And O would be (1*16.00) =16.00

OpenStudy (anonymous):

Am I wrong?

OpenStudy (somy):

yep keep goin u r right

OpenStudy (anonymous):

16.00+2.016 =18.016

OpenStudy (somy):

yep so rounding off it'll be 18

OpenStudy (anonymous):

Yes

OpenStudy (somy):

now use mole formula and find mole

OpenStudy (anonymous):

I kinda confused on what goes where

OpenStudy (somy):

mole = mass/ molecular mass

OpenStudy (anonymous):

thanks!

OpenStudy (somy):

so now you find the moles then multiply what you got by Avogadro constant

OpenStudy (anonymous):

mol= (2.46)/(18) =.136666

OpenStudy (anonymous):

So 8.23 x 10^22

OpenStudy (anonymous):

Or did I do it wrong?

OpenStudy (somy):

you did it right

OpenStudy (anonymous):

Thank you for the help!

OpenStudy (somy):

you are welcome :D